OpenStudy (anonymous):

My homework is asking me to Find an equation of the tangent line at the point (pi/6, y(pi/6), what is y(pi/6)?

ganeshie8 (ganeshie8):

y = f(x) = ?

ganeshie8 (ganeshie8):

where is the rest of the question ?

OpenStudy (anonymous):

y=sin(5x)+cos(4x)

ganeshie8 (ganeshie8):

think of that function as : y = f(x) = sin(5x) + cos(4x)

ganeshie8 (ganeshie8):

y(pi/6) = f(pi/6) = sin(5pi/6) + cos(4pi/6)

ganeshie8 (ganeshie8):

they just mean the value of function evaluated at x=pi/6

OpenStudy (anonymous):

Is x equal to sin(5pi/6) + cos(4pi/6) then?

myininaya (myininaya):

y = sin(5pi/6)+cos(4pi/6) when x=pi/6

OpenStudy (anonymous):

It just clicked with me! Thanks for the help!!!